OpenStudy (anonymous):

log 4 x + log 4 (x - 3) = 1

OpenStudy (muzzack):

x = 1

OpenStudy (anonymous):

\[\log4{x}+\log4{x-3} =1\]

OpenStudy (muzzack):

x/(x+3) = 1/4 4x = x+3 x = 1

OpenStudy (anonymous):

4 is base

OpenStudy (mathmale):

suggestions: You might want to explain in words that these are "logs to the base 4", OR present your equation as a drawing (using the Draw utility, below). Also, please include the instructions with this problem.

OpenStudy (anonymous):

A. x = 6 over 5 B. x = 2 C. x = -3 D. x = 4

OpenStudy (mathmale):

\[ \log 4 x + \log 4 (x - 3) = 1\]

OpenStudy (mathmale):

should be written as\[\log_{4}x+\log_{4}(x-3)=1 \]

OpenStudy (anonymous):

... ok...

OpenStudy (mathmale):

Since \[\log a + \log b = \log (a*b),\]

OpenStudy (anonymous):

sox=4

OpenStudy (anonymous):

or -1

OpenStudy (mathmale):

\log_{4}x+\log_{4}(x-3)=1\[\log_{4}x+\log_{4}(x-3)=1~becomes~\log_{4}x(x-3)=1 \]

OpenStudy (mathmale):

Note that we multiply x and (x-3); we do not divide. Know why?

OpenStudy (anonymous):

I know

OpenStudy (mathmale):

How would you now solve for x?

OpenStudy (anonymous):

is log law

OpenStudy (anonymous):

...x(x+3)=4

OpenStudy (mathmale):

Good. What next?

OpenStudy (anonymous):

x=1?

OpenStudy (mathmale):

You can always check your own work by substituting your solution back into the original equation. If you believe x=1, then substitute that into \[\log_{4}x+\log_{4}(x-3)=1\]

OpenStudy (mathmale):

and determine whether the resulting equation is true or not.

OpenStudy (anonymous):

NVM I get it THANKS

OpenStudy (mathmale):

If x=1, then log to the base 4 of (1-3) is undefined. Therefore, x=1 is not a solution. You're welcome!